# thanks: corresponding author

### v2

@@ I Introduction
The most important property of the Lindblad equation is that it is a completely positive and trace-preserving (CPTP) map [5, 6, 20], which implies that the time evolution generated by $\mathcal{L}$ will keep ${\rm Tr}\rho(t)$ unchanged and will preserve the semi-definiteness of $\rho(t)$ at any future time $t$  [^1] The condition of the dynamics being “completely” positive is more general than simply keeping a non-negative $\rho(t_{0})$ non-negative at future times $t>t_{0}$, which is the positivity condition. In other words, not all positive dynamics are completely positive: see Ref. [20] for more details.. The CPTP conditions ensure that a physical initial state $\rho(t_{0})$ remains physical at all future times $t>t_{0}$, where by a physical (mixed) state we mean a positive semi-definite and Hermitian density matrix with unit trace. Note that the Lindblad equation is a CPTP map for any given choice of Hamiltonian $H$ and jump operators $L_{\mu}$ as long as $\Gamma_{\mu}\geq 0$. Specifically, if all $\Gamma_{\mu}=0$, the Lindblad equation reduces to the von Neumann equation $\dot{\rho}=-i[H,\rho]$ that describes the unitary evolution of an isolated quantum system.
@@
In the presence of a time-independent Hamiltonian and time-independent jump operators $L_{\mu}$, the evolution from a generic initial state typically contains two components: a transient component that eventually decays out, and a non-decaying component that is a linear combination of eigenoperators of $\mathcal{L}$ whose eigenvalues have zero real parts (see Appendix A). As a consequence of the relation between the dynamics of the system density matrix $\rho(t)$ and the eigenoperators of the superoperator $\mathcal{L}$, the spectrum of $\mathcal{L}$ has recently been an object of significant interest [22, 23, 24, 25]. It is worth noting that each individual eigenoperator of $\mathcal{L}$ is generically not a physical density matrix as most of them are non-Hermitian and traceless.
@@
Recently, attention has turned towards the dynamics of periodically driven open quantum systems, where the competition between coherent external driving and incoherent dissipation has been experimentally observed to stabilize non-trivial non-equilibrium steady states (NESS) in e.g., coupled arrays of circuit QED resonators [26, 27]. In such open Floquet systems [28, 29, 30, 31, 32], the Liouville superoperator satisfies $\mathcal{L}(t+T)=\mathcal{L}(t)$ (with $T$ the period). For closed quantum systems, such time-dependent problems are solved using Floquet theory, which generalizes the energy eigenstates of time-independent systems to Floquet eigenmodes with the same period $T$ (see Ref. [33] and Appendix B for details). However, due to the intrinsic non-unitary evolution of open quantum systems, strategies from the Floquet theory for closed systems cannot be directly applied in this setting [34]. For instance, prior studies have shown that the concept of a Floquet Hamiltonian (or more generically, a Floquet Liouville superoperator), which exists for periodically driven closed quantum systems, does not always exist for open quantum systems [32, 35, 36, 37] and the non-unitary evolution can lead to difficulties in properly defining Floquet eigenmodes.
@@ II Floquet-Lindblad eigenmodes
where the Liouville superoperator is time-periodic $\mathcal{L}(t)=\mathcal{L}(t+T)$. In this article, we only consider systems with a finite-dimensional Hilbert space $\mathscr{H}$, with the corresponding Liouville space given by $\mathscr{H}\otimes\mathscr{H}^{*}$. Note that while in realistic contexts the explicit time-dependence is often contained solely in the Hamiltonian (as a result of e.g., a coherent external drive), our discussion encapsulates the general case in which both the Hamiltonian and dissipation pieces can be time-periodic.
@@
where $\varrho_{j}(t_{0}+\tau)\equiv\mathcal{U}(t_{0}+\tau,t_{0})e^{-\lambda_{j}\tau}\rho_{j}$ is constructed such that that $\varrho_{j}(t+T)=\varrho_{j}(t)$ and is called a Floquet-Lindblad eigenmode. Thus, the physical significance of the eigenoperator $\rho_{j}$ is that, if the system begins to evolve from an eigenoperator $\rho_{j}=\varrho_{j}(t_{0})$ at time $t_{0}$, it returns to itself up to a factor $e^{\lambda_{j}T}$ after one period. Amongst all the eigenoperators, those with eigenvalues $|e^{\lambda_{j}T}|=1$ are special as these at most accumulate a phase factor at the end of each period i.e., these are non-decaying eigenmodes and are referred to as “coherent Floquet modes”. Moreover, the eigenoperator(s) $\rho_{0}$ with eigenvalue $e^{\lambda_{0}T}=1$ evolves with period $T$ because $\lambda_{0}=0$ implies $\rho_{0}(t)=\varrho_{0}(t)$ by Eq.(6). We call such periodically evolving states with period $T$ as Floquet-Lindblad steady states. If such an eigenoperator $\rho_{0}$ is physical, its evolution $\rho_{0}(t)$ is a NESS of the system. In general, the NESS for the Floquet-Lindblad equation is defined as any physical density operator $\tilde{\rho}(t)$ that solves the equation and obeys $\mathcal{U}_{F}\tilde{\rho}(t_{0})=\tilde{\rho}(t_{0})$.
@@
    It is important to distinguish between the non-decaying component of an initial state and that of a time-dependent solution. The former is a linear combination of $\rho_{j}$ with eigenvalues of $|e^{\lambda_{j}T}|=1$ and thus lies in the non-decaying subspace, while the latter is a linear combination of the evolved eigenoperators $\rho_{j}(t)$ or eigenmodes $\varrho_{j}(t)$, which is not guaranteed to be in the non-decaying subspace unless at the end of each period.
@@ IV Nonlinear Optical Effects in Solids
First, we write the Hamiltonian of the system that is driven by an external periodic drive. Among the many energy bands in an insulator, we focus on the two bands close to the Fermi surface: a fully occupied valance band and an empty conduction band that is involved in the transition induced by a monochromatic electric field (light): $E(t)=Ee^{-i\Omega t}+E^{*}e^{i\Omega t}$. We apply the Shirley-Floquet formalism [43] (discussed in Appendix D and E) and work within the RWA. Thus, our starting point is a 2-by-2 Shirley-Floquet Hamiltonian matrix at each quasi-momentum $\mathbf{k}$ [38] (see Appendix F for more details):
@@
$\bm{h}_{SF}^{\rm RWA}(\mathbf{k})$ $=\begin{pmatrix}\epsilon_{1}^{0}+\Omega&-iA^{*}v_{12}^{0}\\ iAv_{21}^{0}&\epsilon_{2}^{0}\end{pmatrix}\equiv\epsilon I+\bm{d}\cdot\bm{\sigma}$ (20)
$=\begin{pmatrix}\epsilon+d_{z}&d_{x}-id_{y}\\ d_{x}+id_{y}&\epsilon-d_{z}\end{pmatrix},$
where $\epsilon_{1,2}^{0}$ are $\mathbf{k}$-dependent dispersions in the absence of the electric field, $A=E/\Omega$ is the amplitude of the vector potential, and the $\mathbf{k}$-dependent operator $v^{0}$ represents the coupling between the system and the electromagnetic field as discussed in Appendix F. We will use real parameters $\epsilon$ and $\bm{d}=(d_{x},d_{y},d_{z})$ for simplicity.
@@
Then, we couple the system with a photonic heat bath and write down the matrix equation for solving Floquet-Lindblad eigenmodes using the Shirley-Floquet formalism discussed in Appendix E (we omit all the $\mathbf{k}$ and “RWA” superscripts on the matrices here to ease notations)
$\lambda_{j}\bm{\varrho}_{j}=-i[\bm{h}_{SF},\bm{\varrho}_{j}]+\sum_{\mu=1,2}\Gamma_{\mu}\left(\bm{L}_{\mu}\bm{\varrho}_{j}\bm{L}_{\mu}^{\dagger}-\frac{1}{2}\{\bm{L}_{\mu}^{\dagger}\bm{L}_{\mu},\bm{\varrho}_{j}\}\right),$ (21)
where $\bm{L}_{1}=\sigma_{+}$ and $\bm{L}_{2}=\sigma_{-}$ represent the emission/absorption of energy to/from the heat bath, respectively. As usual, for a heat bath with non-negative temperature, $\Gamma_{1}$ is always larger than $\Gamma_{2}$. Specifically, at zero temperature, $\Gamma_{2}=0$ while $\Gamma_{1}$ is positive.
@@
Next, we solve the Floquet-Lindblad steady state by setting the eigenvalue $\lambda_{j}=0$. To solve Eq. (21), we parameterize the steady-state density matrix as $\bm{\rho}_{ss}^{\rm RWA}(\mathbf{k})=\bm{\varrho}_{0}^{\rm RWA}(\mathbf{k})=\frac{1}{2}(I+\braket{\bm{\sigma}}_{ss}^{\rm RWA}\cdot\bm{\sigma})$, imposing the Hermitian and unit trace conditions. Then Eq. (21) becomes
@@
The above calculation is in the RWA since we only keep the two lowest-frequency components in the entire Fourier series. However, it is not difficult to numerically include more terms in the Shirley-Floquet Hamiltonian to go beyond RWA and get nonlinear optical responses up to any order all at once.
@@ Appendix D Shirley-Floquet Formalism and the Extended Hilbert Space
$\Longrightarrow\left(H(t)-i\frac{d}{dt}\right)$ |$\ket{u_{j}(t)}=\epsilon_{j}\ket{u_{j}(t)}.$
@@ Appendix E Shirley-Floquet formalism for Floquet-Lindblad equations
$\Longrightarrow$ $e^{\lambda_{j}(t-t_{0})}\mathcal{L}(t)\varrho_{j}(t)-\lambda_{j}e^{\lambda_{j}(t-t_{0})}\varrho_{j}(t)-e^{\lambda_{j}(t-t_{0})}\frac{d\varrho_{j}(t)}{dt}=0$
$\Longrightarrow$ $\mathcal{L}(t)\varrho_{j}(t)-\frac{d}{dt}\varrho_{j}(t)=\lambda_{j}\varrho_{j}(t)$
$\Longrightarrow$ $\mathcal{K}(t)\varrho_{j}(t)=\lambda_{j}\varrho_{j}(t).$

### v3

@@ I Introduction
The most important property of the Lindblad equation is that it generates a quantum dynamical map that is completely positive and trace-preserving (CPTP) [5, 6, 20], which implies that the time evolution generated by $\mathcal{L}$ will keep ${\rm Tr}\rho(t)$ unchanged and will preserve the semi-definiteness of $\rho(t)$ at any future time $t$  [^1] The condition of the dynamics being “completely” positive is more general than simply keeping a non-negative $\rho(t_{0})$ non-negative at future times $t>t_{0}$, which is the positivity condition. In other words, not all positive dynamics are completely positive: see Ref. [20] for more details.. The CPTP conditions ensure that a physical initial state $\rho(t_{0})$ remains physical at all future times $t>t_{0}$, where by a physical (mixed) state we mean a positive semi-definite and Hermitian density matrix with unit trace. Note that the dynamical map generated by $\mathcal{L}$ is CPTP for any given choice of Hamiltonian $H$ and jump operators $L_{\mu}$ as long as $\Gamma_{\mu}\geq 0$. Specifically, if all $\Gamma_{\mu}=0$, the Lindblad equation reduces to the von Neumann equation $\dot{\rho}=-i[H,\rho]$ that describes the unitary evolution of an isolated quantum system.
@@
In the presence of a time-independent Hamiltonian and time-independent jump operators $L_{\mu}$, the evolution from a generic initial state typically contains two components: a transient component that eventually decays out, and a non-decaying component that is a linear combination of eigenoperators of $\mathcal{L}$ whose eigenvalues have zero real parts (see Appendix A). As a consequence of the relation between the dynamics of the system density matrix $\rho(t)$ and the eigenoperators of the superoperator $\mathcal{L}$, the spectrum of $\mathcal{L}$ has recently been an object of significant interest [22, 23, 24, 25, 26]. It is worth noting that each individual eigenoperator of $\mathcal{L}$ is generically not a physical density matrix as most of them are non-Hermitian and traceless.
@@
Recently, attention has turned towards the dynamics of periodically driven open quantum systems, where the competition between coherent external driving and incoherent dissipation has been experimentally observed to stabilize non-trivial non-equilibrium steady states (NESS) in e.g., coupled arrays of circuit QED resonators [27, 28]. In such open Floquet systems [29, 30, 31, 32, 33], the Liouville superoperator satisfies $\mathcal{L}(t+T)=\mathcal{L}(t)$ (with $T$ the period). For closed quantum systems, such time-dependent problems are solved using Floquet theory, which generalizes the energy eigenstates of time-independent systems to Floquet eigenmodes with the same period $T$ (see Ref. [34] and Appendix B for details). However, due to the intrinsic non-unitary evolution of open quantum systems, strategies from the Floquet theory for closed systems cannot be directly applied in this setting [35]. For instance, prior studies have shown that the concept of a Floquet Hamiltonian (or more generically, a Floquet Liouville superoperator), which exists for periodically driven closed quantum systems, does not always exist for open quantum systems [33, 36, 37, 38] and the non-unitary evolution can lead to difficulties in properly defining Floquet eigenmodes (see Ref. [39] for an algorithm for getting NESS without involving Floquet formalism).
@@ II Floquet-Lindblad eigenmodes
We focus on the problem of how an open quantum system interacts with a periodically varying time-dependent external field with period $T$. We limit ourselves to the simple situation in which the external field is only coupled to the system, such that the environment remains at thermal equilibrium at all times. Moreover, we assume throughout that we are in a regime in which the Born-Markov approximations hold. Consequently, after integrating out the degrees of freedom of the environment, the equation of motion for the reduced density matrix $\rho(t)$ of the system takes the form Eq. (1), i.e., for each fixed time $t$ the dissipation terms are in Lindblad form, with both the Hamiltonian $H(t)$ and the jump operators $L_{\mu}(t)$ depending explicitly on time $t$ and having the same period $T$ as the external field [4]. As the microscopic derivation of the equation of motion in Lindlbad form involves the Floquet theory, we will refer to it as the Floquet-Lindblad equation [41, 42, 43, 32, 44, 45, 38] throughout this paper.
@@
where the Liouville superoperator is time-periodic $\mathcal{L}(t)=\mathcal{L}(t+T)$. In this article, we only consider systems with a finite-dimensional Hilbert space $\mathscr{H}$, with the corresponding Liouville space given by $\mathscr{H}\otimes\mathscr{H}^{*}$. We emphasize here that the Floquet-Lindblad equation only applies to the situations where the Born-Markov approximation holds [4], otherwise the dynamics of the environment will generically make the system’s dynamics non-local in time and thus cannot be simply expressed by a superoperator $\mathcal{L}(t)$. In more general cases beyond Born-Markov approximation, not only does the master equation fail to exhibit the structure described in Eq. (2), but also a time-periodic external driving does not necessarily result in periodic time dependence.
@@
where $\varrho_{j}(t_{0}+\tau)\equiv\mathcal{U}(t_{0}+\tau,t_{0})e^{-\lambda_{j}\tau}\rho_{j}$ is constructed such that that $\varrho_{j}(t+T)=\varrho_{j}(t)$ and is called a Floquet-Lindblad eigenmode. Thus, the physical significance of the eigenoperator $\rho_{j}$ is that, if the system begins to evolve from an eigenoperator $\rho_{j}=\varrho_{j}(t_{0})$ at time $t_{0}$, it returns to itself up to a factor $e^{\lambda_{j}T}$ after one period. Amongst all the eigenoperators, those with eigenvalues $|e^{\lambda_{j}T}|=1$ are special as these at most accumulate a phase factor at the end of each period i.e., these are non-decaying eigenmodes and are referred to as “coherent Floquet modes”. Moreover, the eigenoperator(s) $\rho_{0}$ with eigenvalue $e^{\lambda_{0}T}=1$ evolves with period $T$ because $\lambda_{0}=0$ implies $\rho_{0}(t)=\varrho_{0}(t)$ by Eq. (6). Such periodically evolving states with period $T$ are referred to as Floquet-Lindblad steady states [44, 45, 38]. If such an eigenoperator $\rho_{0}$ is physical, its evolution $\rho_{0}(t)$ is a NESS of the system. In general, the NESS for the Floquet-Lindblad equation is defined as any physical density operator $\tilde{\rho}(t)$ that solves the equation and obeys $\mathcal{U}_{F}\tilde{\rho}(t_{0})=\tilde{\rho}(t_{0})$.
@@
    It is important to distinguish between the non-decaying component of an initial state and that of a time-dependent solution. The former is a linear combination of $\rho_{j}$ with eigenvalues of $|e^{\lambda_{j}T}|=1$ and thus lies in the non-decaying subspace, while the latter is a linear combination of the evolved eigenoperators $\rho_{j}(t)$ or eigenmodes $\varrho_{j}(t)$, which is not guaranteed to be in the non-decaying subspace unless at the end of each period. The properties of the non-decaying component of an initial state were previously discussed from the perspective of Floquet dynamical symmetries (FDS) in Ref. [45].
@@ IV Nonlinear Optical Effects in Solids
Figure 2: The structure of the effective two-level system at a quasi-momentum $\mathbf{k}$. Indices $1$ and $2$ respectively denote the valence band and the conduction band, each having its own energy dispersion $\epsilon_{1,2}^{0}(\mathbf{k})$. The applied time-periodic light field has angular frequency $\Omega$.
@@
First, we write the Hamiltonian of the system that is driven by an external periodic drive. Among the many energy bands in an insulator, we focus on the two bands close to the Fermi surface: a fully occupied valance band and an empty conduction band that is involved in the transition induced by a monochromatic electric field (light): $E(t)=Ee^{-i\Omega t}+E^{*}e^{i\Omega t}$, see Fig. 2. The many-body Hamiltonian of the solid system in the absence of the driving field and within the subspace of these two energy bands is
$H^{0}=\sum_{\mathbf{k}}\sum_{m=1}^{2}\epsilon_{m}^{0}(\mathbf{k})c_{\mathbf{k},m}^{\dagger}c_{\mathbf{k},m},$ (20)
where $c_{\mathbf{k},m}^{\dagger}$ denotes the creation operator of the electron at quasi-momentum $\mathbf{k}$ and energy band $m$. The band indices $m=1,2$ respectively denote the valence and conduction bands, each of which having its own energy dispersion $\epsilon_{1,2}^{0}(\mathbf{k})$. The coupling between the electrons and the light field introduces a periodically time-dependent term in the Hamiltonian [40]:
$H(t)=H^{0}+A(t)V^{0},$ (21)
where $V^{0}=\sum_{\mathbf{k},m,n}c_{\mathbf{k},m}^{\dagger}v^{0}_{mn}(\mathbf{k})c_{\mathbf{k},n}$ is the current operator in the absence of the driving field (see Appendix F), and $A(t)=-iAe^{-i\Omega t}+iA^{*}e^{i\Omega t}$ is the vector potential corresponding to the driving field: $E(t)=-\frac{\partial A(t)}{\partial t}$, such that its amplitude is $A=E/\Omega$.
Since the form of the Hamiltonian (21) ensures that quasi-momentum $\mathbf{k}$ is conserved during evolution, we can treat each $\mathbf{k}$ sector as an independent two-level single-particle quantum system to be dealt with separately. In the basis spanned by single particle states {|$\{\ket{\mathbf{k},1},\ket{\mathbf{k},2}\}$, the Hamiltonian of the two-level system is represented by a 2-by-2 matrix:
We now apply the Shirley-Floquet formalism [50] (discussed in Appendix D and E), which expands the time dependent Hamiltonian (22) into an infinite dimensional matrix consisting the Fourier coefficients of $\bm{h}(\mathbf{k},t)$. For simplicity of the demonstration, we will work within the RWA by only keeping a 2-by-2 block of the entire Shirley-Floquet Hamiltonian in each sector $\mathbf{k}$ [40] (see Appendix F for more details):
@@
$\bm{h}_{SF}^{\rm RWA}(\mathbf{k})$ $=\begin{pmatrix}\epsilon_{1}^{0}(\mathbf{k})+\Omega&-iA^{*}v_{12}^{0}(\mathbf{k})\\ iAv_{21}^{0}(\mathbf{k})&\epsilon_{2}^{0}(\mathbf{k})\end{pmatrix}\equiv\epsilon I+\bm{d}\cdot\bm{\sigma}$ (23)
$=\begin{pmatrix}\epsilon+d_{z}&d_{x}-id_{y}\\ d_{x}+id_{y}&\epsilon-d_{z}\end{pmatrix}.$
We will use real parameters $\epsilon$ and $\bm{d}=(d_{x},d_{y},d_{z})$ to denote them for simplicity.
@@
Then, we couple the system with a photonic heat bath and write down the matrix equation for solving Floquet-Lindblad eigenmodes using the Shirley-Floquet formalism discussed in Appendix E
$\lambda_{j}\bm{\varrho}_{j}^{\rm RWA}(\mathbf{k})$ $=-i[\bm{h}_{SF}^{\rm RWA}(\mathbf{k}),\bm{\varrho}_{j}^{\rm RWA}(\mathbf{k})]$ (24)
$+\sum_{\mu=1,2}\Gamma_{\mu}(\mathbf{k})$ $\left(\bm{L}_{\mu}\bm{\varrho}_{j}^{\rm RWA}(\mathbf{k})\bm{L}_{\mu}^{\dagger}-\frac{1}{2}\{\bm{L}_{\mu}^{\dagger}\bm{L}_{\mu},\bm{\varrho}_{j}^{\rm RWA}(\mathbf{k})\}\right),$
where $\bm{L}_{1}=\sigma_{+}$ and $\bm{L}_{2}=\sigma_{-}$ represent the emission/absorption of energy to/from the heat bath, respectively. At any quasi-momentum $\mathbf{k}$, the corresponding damping rates $\Gamma_{1,2}$ are controlled by the average photon number $N$ in the heat bath [4]:
$\Gamma_{1}=\gamma_{0}(N+1),\ \ \ \Gamma_{2}=\gamma_{0}N,$ (25)
where $\gamma_{0}$ is a positive coupling constant given by the microscopic interaction between the photons in the bath and the electrons in the solid. The average number $N$ of photons at a certain energy in the bath is determined by the bath temperature via the Planck distribution [4]:
$N=\frac{1}{\exp(\beta\varepsilon)-1},$ (26)
where $\varepsilon=\epsilon_{2}^{0}(\mathbf{k})-\epsilon_{1}^{0}(\mathbf{k})$ is the photon energy and $\beta$ is the inverse temperature of the heat bath. As the bath temperature is encoded in $\Gamma_{1}$ and $\Gamma_{2}$, the Floquet-Lindblad method is capable of discussing the system coupled to heat baths at arbitrary temperatures. In this setup, the emission rate $\Gamma_{1}$ is always larger than the absorption rate $\Gamma_{2}$. Specifically, at zero temperature, $N=0$ so that $\Gamma_{2}=0$ and $\Gamma_{1}=\gamma_{0}$ is positive.
@@
Next, we solve the Floquet-Lindblad steady state by setting the eigenvalue $\lambda_{j}=0$. To solve Eq. (24), we parameterize the steady-state density matrix as $\bm{\rho}_{ss}^{\rm RWA}(\mathbf{k})=\bm{\varrho}_{0}^{\rm RWA}(\mathbf{k})=\frac{1}{2}(I+\braket{\bm{\sigma}}_{ss}^{\rm RWA}\cdot\bm{\sigma})$, imposing the Hermitian and unit trace conditions. Then, (suppressing the $\mathbf{k}$ index) Eq. (24) becomes
@@
The above demonstrative calculation is in the RWA as we only keep the two lowest-frequency components in the entire Fourier series, reducing the Shirley-Floquet Hamiltonian at each momentum sector $\mathbf{k}$ to a 2-by-2 matrix (23). To go beyond RWA in numerical calculations, one can simply include more terms with higher frequencies in the infinite dimensional Shirly-Floquet Hamiltonian matrix, getting nonlinear optical responses up to a higher order all at once.
@@ Appendix D Shirley-Floquet Formalism and the Extended Hilbert Space
and thus
$\left(H(t)-i\frac{d}{dt}\right)\ket{u_{j}(t)}=\epsilon_{j}\ket{u_{j}(t)}.$ (60)
@@ Appendix E Shirley-Floquet formalism for Floquet-Lindblad equations
and hence $\mathcal{K}(t)$ has the same spectrum ${\lambda_{j}}$,
$\mathcal{K}(t)\varrho_{j}(t)=\lambda_{j}\varrho_{j}(t).$ (71)